Status Report

Hubble Space Telescope Daily Report #2916 18 Jul 2001

By SpaceRef Editor
July 18, 2001
Filed under , ,



PERIOD COVERED: 0000Z (UTC) 07/17/01 – 0000Z (UTC) 07/18/01

Daily Status Report as of 199/0000Z


1.1 Completed WF/PC-2 9160 (Disks and Envelopes of Nearby Nebulous
Young Stellar Objects: A Snapshot Survey)

The WF/PC-2 was used to perform a snapshot survey of nearby
nebulous young stellar objects to study the detailed morphology of the
their disks and envelopes and probe the effect of inclination on the
infrared spectral energy distribution of disk/envelope systems. The
proposal completed as planned.

1.2 Completed Eleven Sets of STIS/CCD 9248 (Probing the Large Scale
Structure: Cosmic Shear Observations)

The Space Telescope Imaging Spectrograph (CCD) was used to measure
the distortion of light bundles from distant galaxies that probe the
statistical properties of the intervening inhomogeneous {dark} matter
distribution. Its tidal gravitational field distorts the observable image
shapes thereby causing a coherent ellipticity pattern {Cosmic Shear}. The
observations completed nominally.

1.3 Completed Three Sets of STIS/CCD 8864 (CCD Dark Monitor-Part 2)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the darks for the CCD. The proposal completed nominally.

1.4 Completed Three Sets of WF/PC-2 9255 (Cycle 10 WF/PC-2 Astrometric
Effects of CTE)

The WF/PC-2 was used to quantify the astrometric effects of CTE by
measuring: {1} the relative separation of a bright source vs. a faint
target at different positions on the PC1 CCD and {2} the relative motion of
a source on the CCD compared to very precise slews performed with the
FGSs. These tests will be conducted for point and extended targets at
several different intensity levels. No problems were encountered.

1.5 Completed Six Sets of WF/PC-2 9252 (Cycle 10 Clocks On Check)

The WF/PC-2 was used to perform closure calibration for clocks ‘on’
mode, which obtains dark frames in order to provide data to create a clocks
‘on’ super dark calibration reference file, to monitor and characterize the
evolution of hot pixels, and to obtain standard star observations for
photometric calibration. The proposal completed with no reported anomalous

1.6 Completed STIS/CCD 8865 (Bias Monitor-Part 2)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at
gain = 4, to build up high-S/N superbiases and track the evolution of hot
columns. The proposal completed with no anomalous activity.

1.7 Completed WF/PC-2 9085 (Measuring The Masses Of High-Z Quasar Host

The WF/PC-2 was used to investigate the close link between
black-hole mass and spheroid mass in both quiescent and active galaxies in
the local universe. Consequently black-hole and spheroid formation/growth
are now viewed as intimately related processes, and establishing the mass
of quasar host galaxies as a function of redshift is now seen as a key
measurement in observational cosmology. From the Cycle 7 NICMOS program,
the best estimate to date of the mass evolution of the hosts of both
radio-quiet and radio-loud quasars out to z ~eq 2 has been derived. Under
the assumption of passive stellar-population evolution our results are
consistent with the black-hole/spheroid population being unchanged out to z
~eq 2. However, the crucial assumption of passive evolution needs to be
tested. All observations competed nominally.

1.8 Completed Three Sets of WF/PC-2 8828 (Cycle 9 Supplemental Darks pt3/3)

The WF/PC-2 was used to obtain three dark frames every day to
provide data for monitoring and characterizing the evolution of hot
pixels. The proposal completed with no reported problems.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 6

Successful: 6

Scheduled Re-acquisitions: 10

Successful: 10

2.2 FHST Updates:

Scheduled: 10

Successful: 10

2.3 Operations Notes:

The SSR EDAC error counter was cleared twice per ROP SR-1A.

As described in HSTAR 8276, the MCEA electronics temperature
parameter (SMAELEXT) flagged out-of-limits high with a value of 49.0925 at
198/175633Z. The upper limit is 49. This temperature sensor is in bay 8,
toward the +V3 axis, facing the Sun.

The STIS MCE-1 reset at 198/092055Z, while the high voltage was off
and while inside an SAA interval. MAMA-1 will be recovered next week via
normal SMS commanding, occurring at the time of the next high voltage on
commanding. The STIS flight software error counter was reset at 198/1002Z,
using ROP NS-12.


Continuation of normal science observations and calibrations.

SpaceRef staff editor.